Find the Partial Fraction Decompostion for -13x-37/x4-x3+x2-3x-6
Answer #1

factoring x^4 - x^3 + x^2 - 3x - 6

= ( x + 1 ) ( x- 2) (x^2 + 3)

so, we can write

( - 13x - 37 ) / ( x + 1 ) ( x- 2) (x^2 + 3) = ao / ( x+ 1) + a1 / ( x - 2) + ( a3x + a2 ) / ( x^2 + 3)

( - 13x - 37 ) / ( x + 1 ) ( x- 2) (x^2 + 3) = ao ( x - 2) ( x^2 + 3) + a1 ( x +1) ( x^2 + 3) + ( a3x + a2 ) ( x+ 1 )( x - 2) / ( x + 1 ) ( x- 2) (x^2 + 3)

a0 = 1 , a 1 = - 3

a3 = 1 , a2 = 8

hence, partial fraction decomposition is